# Prediction-Enhanced Monte Carlo: A Machine Learning View on Control Variate

### v2

@@ Abstract.
For many complex simulation tasks spanning areas such as healthcare, engineering, and finance, Monte Carlo (MC) methods are invaluable due to their unbiased estimates and precise error quantification. Nevertheless, Monte Carlo simulations often become computationally prohibitive, especially for nested, multi-level, or path-dependent evaluations lacking effective variance reduction techniques. While machine learning (ML) surrogates appear as natural alternatives, naïve replacements typically introduce unquantifiable biases. We address this challenge by introducing Prediction-Enhanced Monte Carlo (PEMC), a framework that leverages modern ML models as learned predictors, using cheap and parallelizable simulation as features, to output unbiased evaluation with reduced variance and runtime. PEMC can also be viewed as a “modernized" view of control variates, where we consider the overall computation-cost-aware variance reduction instead of per-replication reduction, while bypassing the closed-form mean function requirement and maintaining the advantageous unbiasedness and uncertainty quantifiability of Monte Carlo.
We illustrate PEMC’s broad efficacy and versatility through three examples: first, equity derivatives such as variance swaps under stochastic local volatility models; second, interest rate derivatives such as swaption pricing under the Heath-Jarrow-Morton (HJM) interest-rate model.Finally, we showcase PEMC in a socially significant context—ambulance dispatch and hospital load balancing—where accurate mortality rate estimates are key for ethically sensitive decision-making.
Across these diverse scenarios, PEMC consistently reduces variance while preserving unbiasedness, highlighting its potential as a powerful enhancement to standard Monte Carlo baselines.
Equal Contribution. 1Morgan Stanley. 2Department of IEOR, Columbia University. The Work was done during the internship at Morgan Stanley in the Machine Learning Research group.
     Corresponding author: fengpei.li@morganstanley.com
@@ 1. Introduction
Monte Carlo (MC) is a fundamental tool for evaluating complex stochastic models where closed-form solutions are unavailable. Its main strengths—unbiasedness and rigorous uncertainty quantification—make it essential in high-stakes, high-complexity applications like healthcare resource allocation and financial risk management, where fairness or regulatory compliance are crucial. For example, transparent and bias-aware methods in healthcare are vital to ensure equitable outcomes and prevent disparities in care access [110, 76, 41]. Similarly, Monte Carlo estimation is widely used in financial risk management to implement hedging strategies, rebalance portfolios, and establish reliable error bounds, as even small biases in financial simulations can accumulate over time and lead to significant errors [30].
@@
Despite these advantages, classical MC faces significant challenges in nested, multilevel, or path-dependent simulations that are difficult to parallelize. Complex stochastic dynamics can make each sample path extremely costly to generate—sometimes taking hours for one sample (e.g., credit valuation adjustment [62])—and effective variance reduction methods, such as control variates, are often unavailable [9]. Consequently, MC’s inherent $\mathcal{O}(1/\sqrt{n})$ convergence rate often prohibits real-time decision-making with adequate precision. This is also problematic in exotic option pricing, where sample paths are generated sequentially under complex dynamics, making parallelization unavailable [60]. The issue is further exacerbated when many evaluation tasks need to be repeatedly conducted in a timely manner. In application domains such as option pricing in computational finance, such a computational barrier has been pointed out prominently, e.g., [73, Chapter 21] or [30, Chapter 4].
@@
Conversely, machine learning (ML) based predictive modeling has emerged as a fast alternative for approximating MC evaluations, including pricing financial derivatives [74, 52, 15]. Neural networks and other ML techniques can learn complex mappings and, once trained, can generate predictions at negligible marginal cost - what practitioners often call “fast evaluation" [72]. However, the black-box nature and the lack of inherent statistical control lead to unquantifiable biases, rendering them unsuitable for tasks requiring rigorous error quantification, e.g., high-stakes applications in risk management [74, 32]. Thus, MC and ML represent opposite ends of the spectrum: MC is reliable and uncertainty quantifiable but intrinsically slow, while ML is computationally fast but falls short in statistical guarantees.
@@
Given this trade-off between the speed of ML and the reliability of MC, a natural question arises: can we design a method that combines both advantages while avoiding their respective pitfalls? In this paper, we answer affirmatively by introducing the Prediction-Enhanced Monte Carlo (PEMC) framework. PEMC leverages predictive ML modeling to replace a large portion of costly simulations with cheap/parallelizable samples from ML models, reducing variance on a scheme-wide level. Moreover, despite relying on black-box predictive models, PEMC maintains unbiasedness, balancing computational savings with rigorous statistical guarantees.
1.1. Contribution
@@
The PEMC innovates in two aspects. First, PEMC integrates ML predictive models into any MC baseline while preserving unbiasedness and error quantification of MC. Second, we show how the predictive modeling of ML combined with cheap, parallelizable samples will result in a scheme-wide reduction of variance and costly samples, enhancing real-time efficiency for complex MC simulations. PEMC method can also be viewed as a modernized variant of the control variate (CV) technique, a classical approach that reduces Monte Carlo variance via auxiliary outputs whose known mean offsets random fluctuations [60, 9]. Unfortunately, the classical CV requirement of a closed-form mean of the a correlated auxiliary output easily becomes prohibitive in many modern applications or complicated stochastic systems. In such cases, there simply is no readily available analytic expression for the auxiliary variable’s expectation, effectively barring the use of a CV.
@@
The PEMC framework bridges the above gap and substantially enlarge the scope of CV to many ML candidates. The key idea is the following: In the traditional CV framework, the known mean of CV is used to achieve a per-replication variance reduction while maintaining unbiasedness, i.e., the variance is reduced even if we run just one simulation replication. This is a powerful ability, but as mentioned above it is achieved at the expense of severely narrowing the applicability of CV by requiring closed-form mean function. However, if the CV itself is computationally cheap in the sense of being estimatable with fast simulation, then we can afford to run plenty of additional MC separately from the original MC to estimate its mean. This would preserve unbiasedness and overall estimation variance relative to the total computation cost, we could achieve a scheme-wide efficiency improvement. That is, in the PEMC framework, we consider scheme-wide instead of per-replication variance reduction, through which we relax the crucial requirement of known CV mean, and this in turn allows us to substantially expand the CV methodology to cover modern ML models. The alignment of our framework with the use of ML predictions prompts us to regard PEMC as a modernized view of CV.
We demonstrate PEMC’s efficacy across multiple MC applications. First, we consider two exotic financial derivative-pricing scenarios: (i) variance swap pricing under a stochastic local volatility model, a fundamental instrument for volatility trading due to its payout on realized variance; and (ii) swaption pricing within the Heath-Jarrow-Morton (HJM) interest rate framework, a sophisticated yet widely used derivative by financial institutions for hedging and speculation. Both instruments are highly liquid, exchange-traded products but pose significant computational challenges due to the infeasibility of traditional variance reduction techniques, as they lack analytic or easily computable expected values. Additionally, we illustrate PEMC’s broad applicability through an ambulance dispatch and hospital load-balancing simulation, where unbiased mortality rate estimates critically inform ethical and efficient resource allocation. These examples collectively highlight PEMC’s capability to leverage machine learning predictions to effectively manage computational complexity while maintaining unbiased estimation and reliable error quantification.
@@ 1.2. Literature Review
The literature on enhancing Monte Carlo (MC) simulations with machine learning (ML) approximators is extensive, yet many existing ML-based methods for control variates (CV) and variance reduction impose restrictive conditions that limit their practical applicability. Approaches such as reproducing Stein kernels [100, 88], Neural CV derived from Stein’s identity [129, 82], regularized least squares for CV construction [106, 121, 85], adaptive CV schemes [66, 67, 79], and $\mathscr{L}^{2}$ function approximation frameworks [93] have demonstrated theoretical promise with potentially rapid convergence rates. Recent theoretical work by [18] provides important insights into when regression-adjusted control variates can achieve optimal performance, showing that their effectiveness depends critically on function smoothness. However, these methods typically require knowledge of the control variate mean or impose specific structural assumptions that can be difficult to check. As a result, while these existing methods offer strong theoretical foundations, their applicability in complex financial modeling scenarios can be sometimes limited. Multilevel Monte Carlo (MLMC) [58], another prominent variance reduction approach, has shown effectiveness through coupling simulations at different discretization levels but is often challenging to implement, requiring careful design for stable coupling [107, 19, 112, 57, 115]. Our Prediction-Enhanced Monte Carlo (PEMC) framework addresses these limitations comprehensively by eliminating the necessity of known CV means, enabling the direct use of modern ML predictors as control variates, while rigorously preserving unbiasedness.
In a different line of literature, the work of Emsermann and Simon [49] on quasi control variates and the work of [102] on control variates using Estimated Means (CVEMs) share notable similarities with our approach, as it relies on an estimated mean (rather than a priori known, and thus not a traditional control variate) of a quantity that is expected to exhibit a strong correlation with the random variable of interest. However, our methodology distinguishes itself in several critical ways. First, our primary motivation centers on integrating modern ML capabilities to train a suitable control variate, rather than relying on a fixed quantity observed to be correlated with the target (which may not always be obvious or even available, especially for our options pricing examples). In other words, we can create an effective control variate from an easily estimable quantity, even if its naive form does not exhibit strong correlation with the target. This approach circumvents the challenges of identifying a closed-form mean and strictly relying on high correlation-based pairing. Second, instead of adopting an adaptive algorithm as they do, we focus on a pre-training approach, demonstrating that our performance advantages would not emerge under their adaptive scheme. Third, our analysis is not anchored exclusively in asymptotic arguments. We provide a formulation that can be interpreted in both finite-sample and asymptotic regimes, offering more direct and flexible theoretical guarantees.
In contrast, the Prediction-Enhanced Monte Carlo (PEMC) framework overcomes these constraints by reimagining the control variate paradigm to function effectively even when auxiliary means are unknown. Unlike previously discussed methods, PEMC aligns closely with the recently proposed Prediction-Powered Inference (PPI) approach [6], particularly in its hybrid methodology that integrates classical, unbiased statistical principles with modern machine learning (ML) techniques. This hybrid strategy is becoming increasingly important in evaluating generative models such as large language models (LLMs), where rapid, interpretable, and unbiased metrics are essential [135, 22, 51]. PPI-type approaches combine sparse human annotations with abundant ML-generated labels, facilitating accurate model performance evaluation even in low-label scenarios.
Additionally, these methodologies intersect with a rich literature in causal inference, specifically doubly robust estimators [10, 55, 113] and orthogonal statistical learning [36, 54]. Doubly robust methods construct additional robustness layers by pairing sophisticated predictive models with auxiliary estimators, ensuring that minor errors in one model component do not significantly bias the overall estimation. This philosophy closely mirrors the logic of control variates, and by extension, the core principles of our PEMC framework.
Traditional variance reduction and numerical approximation techniques also have a long and established history in computational finance and stochastic modeling. One example conceptually related to PEMC is the Multilevel Monte Carlo (MLMC) method, as introduced and popularized by [58]. MLMC exploits a hierarchy of discretizations (e.g., finer and coarser time steps in an SDE simulation) to reduce variance at a fraction of the cost of a naive Monte Carlo approach. Its basic premise involves coupling simulations at different resolution levels so that differences between levels have lower variance, ultimately leading to significant computational savings. Despite its solid theoretical foundations, MLMC can be challenging to implement in practice: constructing suitable couplings between the fine and coarse approximations, ensuring stability, and tuning parameters to achieve the promised complexity gains are all nontrivial tasks. Nevertheless, the class of MLMC methods and their adoption of a coupling perspective provide extremely valuable insights into how PEMC manages the relationship between the target and the predictor [107, 19, 112, 57, 115].
@@
In computational finance, the emergence of machine learning (ML) has also sparked significant research efforts aimed at enhancing the speed and accuracy of financial computations, including option pricing. By using techniques such as neural network surrogates, these methods can approximate pricing functions or the Greeks at a fraction of the computational cost of traditional numerical schemes [74, 32, 119, 114, 84]. Early work by hutchinson1994nonparametric [74], for instance, used neural networks to relax stringent modeling assumptions, while more recent approaches [32, 119, 114, 84] demonstrate that deep architectures can effectively handle high-dimensional and path-dependent pricing problems. However, these purely data-driven approximations typically lack the rigorous error controls afforded by more classical methods. As a result, biases introduced during training and potential overfitting can lead to systematic mispricings, particularly under market regimes not well-represented in the training data. Such issues underscore the need for frameworks that combine the computational advantages of ML with robust theoretical guarantees to ensure reliable and unbiased pricing outcomes. On the other hand, within option pricing, even though both MC and PDE-based methods are widely used, it is known that PDE methods generally cannot be used to price options (i.e., payoff depends on the price path). While special cases allow for PDE formulations through state-space augmentation or boundary condition adjustments, these remain as exceptions rather than the rule [116].
@@
The rest of the paper is organized as follows. In Section 3, we provide a comprehensive background on PEMC. In Section 4, we provide a detailed and comprehensive analysis of the PEMC estimator, presenting both theoretical results and practical illustrations. In Section 5, we apply PEMC it to real-world exotic option pricing problems of variance swaps pricing under stochastic (local) volatility models and the pricing of swaptions under the HJM framework. We conclude the paper in Section 6 with discussion on broader implications and potential directions for future research.
2. Prediction-Enhanced Monte Carlo
@@
for arbitrary, potentially high-dimensional parameter $\boldsymbol{\theta}\in\boldsymbol{\Theta}\subseteq\mathbb{R}^{d_{\theta}}$, where $\Theta$ is the parameter space. Here, $\boldsymbol{Y}\in\mathbb{R}^{d_{y}}$ represents a high-dimensional random vector governed by $\mathbb{E}_{\boldsymbol{\theta}}$ and $f_{\boldsymbol{\theta}}$ is the evaluation function $\mathbb{R}^{d_{y}}\to\mathbb{R}$. In general, both $f_{\boldsymbol{\theta}}$ and $\mathbb{E}_{\boldsymbol{\theta}}$ can depend on $\boldsymbol{\theta}$. The goal is to produce an unbiased estimate of $\mu(\boldsymbol{\theta})$ for any $\boldsymbol{\theta}\in\boldsymbol{\Theta}$, while the primary bottleneck stems from costly and time-consuming simulation of $\boldsymbol{Y}$.

### v3

@@ Abstract.
For many complex simulation tasks spanning areas such as healthcare, engineering, and finance, Monte Carlo (MC) methods are invaluable due to their unbiased estimates and precise error quantification. Nevertheless, MC simulations often become computationally prohibitive, especially for nested, multi-level, or path-dependent evaluations lacking effective variance reduction techniques. While machine learning (ML) surrogates appear as natural alternatives, naïve replacements typically introduce unquantifiable biases. We address this challenge by introducing Prediction-Enhanced Monte Carlo (PEMC), a framework that leverages modern ML models as learned predictors, using cheap and parallelizable simulations as features, to output unbiased evaluations with reduced variance and runtime. As a result, PEMC eliminates the closed-form-mean requirement that constrains classical control-variate methods, preserves unbiasedness and explicit confidence intervals of MC, and achieves scheme-wide variance reduction. Our theoretical analysis quantifies the optimal allocation between expensive evaluations and large batches of cheap, parallelizable feature draws. Across three representative applications—variance-swap pricing under stochastic-local-volatility, swaption pricing under Heath–Jarrow–Morton models, and ambulance diversion policy evaluation—we show that PEMC reduces root-mean-squared error by 30–55% relative to standard MC at similar computational cost.
Equal Contribution. 1Morgan Stanley. 2Department of IEOR, Columbia University. The work was done during empolyment at Morgan Stanley.
    Corresponding author: fl2412@columbia.edu
@@ 1. Introduction
Monte Carlo (MC) is a fundamental tool for evaluating complex stochastic models where closed-form solutions are unavailable. Its main strengths—unbiasedness and rigorous uncertainty quantification through variance estimation and confidence intervals—make it essential in high-stakes, high-complexity applications such as healthcare resource allocation and financial risk management, where fairness or regulatory compliance are crucial. For example, transparent and bias-aware methods in healthcare are vital to ensure equitable outcomes and prevent disparities in care access [113, 77, 42]. Similarly, financial risk assessments rely on MC estimation to support hedging, re-balancing decisions, and valid error bounds, as minor biases in financial simulations can accumulate [30].
@@
Despite these advantages, classical MC faces significant challenges in nested, multilevel, or path-dependent simulations that are difficult to parallelize. Complex stochastic dynamics can make each sample path extremely costly to generate—sometimes taking hours for one sample (e.g., to estimate credit valuation adjustments in over-the-counter contracts, see [63])—and effective variance reduction methods are often unavailable [9]. Consequently, MC’s inherent $\mathcal{O}(1/\sqrt{n})$ convergence rate often prohibits real-time decision-making with adequate precision. This is the case, for example, in the context of pricing of exotic options, where sample paths need to be generated sequentially under complex dynamics, making parallelization not easily achievable [61]. These computational challenges are critical in financial derivative pricing, as highlighted in the literature, for example, in [74, Chapter 21] and [30, Chapter 4]. Similar computational bottlenecks arise in healthcare simulation, where modeling complex patient flows and hospital operations can require extensive computation time but transparent and bias-aware methods are crucial to ensure equitable outcomes [113, 77, 42].
@@
Conversely, machine learning (ML) based predictive modeling has emerged as a fast alternative for approximating MC evaluations, including pricing financial derivatives [75, 53, 15]. Neural networks and other ML techniques can learn complex mappings and, once trained, can generate predictions at negligible marginal cost - what practitioners often call “fast evaluation" [73]. However, their black-box nature and lack of inherent statistical control hinder the quantification of errors, rendering them unsuitable for tasks requiring rigorous reliability assessments, such as high-stakes applications in risk management [75, 32]. Thus, MC and ML represent opposite ends of the spectrum: MC is reliable and uncertainty quantifiable but intrinsically slow, while ML is computationally fast but falls short in statistical guarantees.
@@
Given this trade-off between the speed of ML and the reliability of MC, a natural question arises: can we design a method that combines both advantages while avoiding their respective pitfalls? In this paper, we provide a positive answer by introducing the Prediction-Enhanced Monte Carlo (PEMC) framework. PEMC leverages predictive ML modeling to replace a large portion of costly simulations with cheap/parallelizable samples from ML models, reducing variance on a scheme-wide level. Moreover, despite relying on black-box predictive models, PEMC maintains unbiasedness, balancing computational savings with rigorous statistical guarantees.
1.1. Our Contributions
@@
PEMC innovates in two aspects. First, PEMC integrates ML predictive models into any MC baseline while preserving unbiasedness and error quantification of MC. Specifically, we show how the predictive modeling of ML combined with cheap, parallelizable samples will result in a reduction of variance and costly samples, enhancing real-time efficiency for complex MC simulations. Second, PEMC can also be viewed as a modernized variant of the control variate (CV) technique, a classical approach that reduces MC variance via auxiliary outputs whose known mean offsets random fluctuations [61, 9, 98]. Unfortunately, the classical CV approach, which requires a closed-form expression for the mean of a correlated auxiliary output, often becomes unavailable in many modern applications or complex stochastic systems, effectively barring the use of CV.
@@
To explain concretely the second point above, our PEMC framework bridges ML with MC by substantially enlarging the scope of CV to many ML candidates. The key idea is the following: In the traditional CV framework, the known mean of CV is used to achieve a per-replication variance reduction while maintaining unbiasedness, i.e., the variance is reduced even if we run just one simulation replication. While this is a powerful feature, it comes at the cost of significantly limiting CV’s applicability, due to the requirement of a closed-form expression for the mean. However, if the CV itself is computationally cheap in the sense of being estimatable with fast simulation, then we can afford to run plenty of additional MC separately from the original MC to estimate its mean. This would preserve unbiasedness and overall estimation variance relative to the total computation cost, as long as we consider efficiency improvement at the scheme-wide level. That is, in the PEMC framework, we consider total cost-aware instead of per-replication variance reduction, through which we relax the crucial requirement of known CV mean, and this in turn allows us to substantially expand the CV methodology to cover modern ML models. The alignment of our framework with the use of ML predictions prompts us to regard PEMC as a modernized view of CV.
@@ 1.2. Literature Review
The literature on enhancing MC simulations with ML approximators is extensive, yet many existing ML-based methods for CV and variance reduction impose restrictive conditions that appear to limit their practical applicability. Approaches such as reproducing Stein kernels [102, 89], Neural CV derived from Stein’s identity [132, 83], regularized least squares for CV construction [109, 124, 86], adaptive CV schemes [67, 68, 80], and $\mathscr{L}^{2}$ function approximation frameworks [94] have demonstrated theoretical promise with potentially rapid convergence rates. Recent theoretical work by [18] provides important insights into when regression-adjusted CVs can achieve optimal performance, showing that their effectiveness depends critically on function smoothness. However, these methods typically require knowledge of the CV mean, in turn restricting the classes of usable CV, or impose specific structural assumptions that can be difficult to check. As a result, while these existing methods offer strong theoretical foundations, their applicability in complex modeling scenarios can be undermined.
In a different line of literature, the works of [50] on quasi CVs and [104] on CVs using Estimated Means (CVEMs) is related to our approach, because they rely on an estimated mean (rather than a priori known, and thus not a traditional CV) of a quantity that is expected to exhibit a strong correlation with the random variable of interest. There are noticeable differences between our methodology and theirs. To begin with, our primary objective is to integrate modern ML capabilities into MC, so that our estimates can retain the error-quantifiability of the latter approach while leveraging the strengths of ML. To this end, a core component of our methodology regards what is needed, and how to train a suitable CV, rather than relying on a prefixed CV candidate, which may not always be obvious or even available. Moreover, our investigation includes the use of a pre-training approach in place of the adaptive algorithms employed in previous works, and provides a framework applicable to both finite-sample and asymptotic regimes, unlike the exclusively asymptotic focus of prior studies.
Within the literature of MC variance reduction, PEMC also relates to Multilevel Monte Carlo (MLMC) methods, originally introduced and popularized by [59], and also the idea of multi-fidelity modeling popular in scientific computation [106]. MLMC reduces variance by leveraging a hierarchy of discretizations (e.g., coarser and finer time steps in SDE simulations) and coupling them to achieve computational efficiency over naive Monte Carlo approaches [58]. The key idea is to couple simulations across resolution levels so that their differences exhibit reduced variance—or even enable unbiased estimators [115, 19, 118]. Along a similar vein, in multi-fidelity modeling, computational simulation models of different levels of resolution, which in turn bring in different accuracies and computational demands, are jointly utilized and concatenated, with input model parameters treated as random variables that allow coupling between the model levels [106]. Note that, despite the strong theoretical foundations, MLMC can be difficult to implement in practice, since constructing effective couplings, ensuring numerical stability, and tuning parameters to realize the theoretical complexity benefits are nontrivial tasks [110]. PEMC adopts the coupling paradigm of MLMC and multi-fidelity modeling but innovates by using ML predictors as CV.
Outside the MC literature, our PEMC framework aligns most closely with the recently proposed Prediction-Powered Inference (PPI) approach [6]. PPI is a statistical framework that enables researchers to construct valid confidence intervals and p-values when combining a small dataset with gold-standard labels and a large dataset with ML predictions. The key challenge PPI addresses is how to leverage abundant but potentially biased ML predictions alongside scarce but trustworthy ground-truth data for statistical inference. Such PPI-type hybrid strategy is becoming increasingly important in evaluating generative models such as large language models (LLMs) [138, 22, 52], where sparse human annotations is combined with abundant ML-generated labels to facilitate accurate model performance evaluation even in low-label scenarios. Additionally, the PPI methodologies also intersect with a rich literature in causal inference, specifically in doubly robust estimators [10, 56, 116] and orthogonal statistical learning [36, 55]. Doubly robust methods construct additional robustness layers by pairing sophisticated predictive models with auxiliary estimators, ensuring that minor errors in one model component do not significantly bias the overall estimation. This philosophy closely mirrors the logic of CV, and by extension, the core principles of our PEMC framework. While PPI replaces expensive gold-standard labels with cheaper ML predictions and auto-evaluation, PEMC focuses on replacing expensive, path-dependent and non-parallelizable MC samples with cheap, efficiently simulated, and highly parallelizable MC samples.
@@
Finally, our work is also related to a branch of literature in computational finance, which has focused on harnessing ML to enhance the speed and accuracy of derivatives pricing. By using techniques such as neural network surrogates, these methods can approximate pricing functions or the Greeks at a fraction of the computational cost of traditional numerical schemes [75, 32, 122, 117, 85]. Early work by hutchinson1994nonparametric [75], for instance, use neural networks to relax stringent modeling assumptions, while more recent approaches [32, 122, 117, 85] demonstrate that deep architectures can effectively handle high-dimensional and path-dependent pricing problems. However, these purely data-driven approximations typically lack the rigorous error controls afforded by more classical methods. In the context of option pricing, even though both MC and partial differential equation (PDE)-based methods are widely used, it is known that PDE methods generally cannot be used to price options (i.e., payoff depends on the price path). While special cases allow for PDE formulations through state-space augmentation or boundary condition adjustments, these remain as exceptions rather than the rule [119].
@@
The rest of the paper is organized as follows. In Section 2, we motivate and introduce PEMC. In Section 3, we provide more details on our framework via a running example of Asian option pricing. In Section 4, we present theoretical results and their practical implications. In Section 5, we apply PEMC to production-grade exotic option pricing, including variance swaps pricing under stochastic (local) volatility models and the pricing of swaptions under the HJM framework, as well as an ambulance diversion problem. We conclude the paper in Section 6 with discussions on broader implications and potential directions for future research.
2. The Prediction-Enhanced Monte Carlo Framework
@@
for arbitrary, potentially high-dimensional parameter $\boldsymbol{\theta}\in\boldsymbol{\Theta}\subseteq\mathbb{R}^{d_{\theta}}$, where $\Theta$ is the parameter space. Here, $\boldsymbol{Y}\in\mathbb{R}^{d_{y}}$ represents a high-dimensional random vector whose probability distribution is parameterized by $\mathbb{E}_{\boldsymbol{\theta}}$ and $f_{\boldsymbol{\theta}}$ is the evaluation function $\mathbb{R}^{d_{y}}\to\mathbb{R}$. In general, both $f_{\boldsymbol{\theta}}$ and $\mathbb{E}_{\boldsymbol{\theta}}$ can depend on $\boldsymbol{\theta}$. The goal is to produce an unbiased estimate of $\mu(\boldsymbol{\theta})$ for any $\boldsymbol{\theta}\in\boldsymbol{\Theta}$, and the primary bottleneck stems from costly simulation of $\boldsymbol{Y}$.
2.1. Existing Challenges of MC, ML, and CV